# Long-range Entanglement and Role of Realistic Interaction in Braiding of Non-Abelian Quasiholes in Fractional Quantum Hall Phases

### v2

@@ III Simultaneous selection of fusion channels
Parity conservation places a constrain on the possible outcome of a fusion of two MR quasiholes. To take a simplest example, let us consider the four-quasihole MR state in the even sector. If one pair of quasiholes is fused with a pre-determined outcome, let’s say, into a $1$-anyon quasihole, then the fusion out come of the other pair is also known. Now, interesting physics can happen due to the following observation: the fusion outcome of two quasiholes directly affect their mutual statistics[48, 57, 56]. This means that by pre-determining the fusion outcome of one pair of quasiholes, one can directly influence the statistics of other pairs of quasiholes in the system regardless of the distance between them. Thus, parity conservation acts like a long-range entanglement between pairs of MR quasiholes (see Fig. 1).
In this section, we explore the physics resulting from this “parity-mediated entanglement”. We start this section by a simple analogy between parity conservation in the MR state and typical picture of entanglement in quantum mechanics (i.e. the Bell state). This will motivate discussions on the effect of spontaneous fusion channel selection on braiding measurement, which we discuss subsequently. Numerical evidences for this effect will also be presented in the next section.
@@ III.1 Parity conservation as long-range entanglement
where $S_{i}\uparrow$ denotes the outcome of measuring spin up in the $i$-th electron. Entanglement in this case is evidence in the fact that “$S_{1}\uparrow$” and “$S_{2}\uparrow$” are not independent events.
Coming back to the MR system, suppose we have four MR quasiholes that are separated into two pairs. Each pair, on its own, can fuse into either $1$ or $\psi$. Due to parity conservation, however, the two pairs must both fuse into $1$ or into $\psi$ (here we take the even sector as an example), so the two possible outcomes can be written as $|\psi,\psi\rangle$ or $|1,1\rangle$. One can then construct a state similar to Eq.(18) as:
@@
Similar to what happens in the Bell state, here the fusion outcome of one pair of anyons, if known, uniquely determines the fusion outcome of the other regardless of their spatial separation. Furthermore, we can also see that Eq.(21) can only exists as the zero-energy state of the model Hamiltonian. Under some two-body interaction, Eq.(21) collapses into either of each component depending on whether $1$-anyon or $\psi$-anyon are energetically favorable. The exact details of the energetics of different anyon species will be discussed in the following section. In this section, we focus on the effect of fusion channel selection of one pair of quasihole on the braiding statistics of the other, and argue that between the two pairs there exists a long-range entanglement mediated by parity conservation.
@@ III.2 Effects on braiding statistics
In a FQH system there are two separate ways the fusion outcome of two anyons can be realized and measured: by placing two anyons at the same spatial co-ordinate, effectively “fusing” them in real space, or by measuring their mutual statistics via braiding. Assuming that during the braiding process the separation between every pair of $\sigma$-anyons remains very large (compared to their sizes), there exists a qualitative difference between the two methods. Namely, the presence of two-body interaction does not affect the degeneracy of the system during the braiding process, but two-body interaction does split the degeneracy when two $\sigma$-anyons are fused in real space. In the latter case, the fusion channel is chosen energetically depending on the details of two-body interaction (see Fig.4).
@@
which resembles Eq. (1) where $\hat{V}_{3}^{3bdy}$ is the three-body interaction the nullspace of which is the Moore-Read quasihole manifold, $\hat{V}_{\psi}^{\text{model}}$ is some “model” two-body interaction that energetically favors $\psi$-anyon over $1$-anyon, and $\hat{V}_{1}^{\text{model}}$ is some other “model” two-body interaction that favors $1$-anyon over $\psi$-anyon, and $\hat{V}_{\text{pins}}$ is a collection of some local one-body pinning potentials. In the next section we shall see that indeed $\hat{V}_{1}^{\text{model}}$ and $\hat{V}_{\psi}^{\text{model}}$ exist, but their exact forms are not important for the current discussion [^3] Note that despite the word “model” used, we do not require that $1$-anyon is the zero-energy state of $\hat{V}_{1}^{\text{model}}$, but only that the $1$-anyon has lower energy than $\psi$-anyon with respect to this interaction. The same is true for the $\psi$-anyon and $\hat{V}_{\psi}^{\text{model}}$.
@@
Using this Hamiltonian, one can design an experiment to observe the effect of long-range entanglement as follows. Suppose, in a MR system with four quasiholes, $\sigma_{1}$,$\sigma_{2}$,$\sigma_{3}$,and $\sigma_{4}$, we pick one pair of quasiholes (say, $\sigma_{1}$ and $\sigma_{2}$) and move their positions to the same spatial point so that they fuse into a charge-$e/2$ quasihole. Afterward, we measure the braiding of $\sigma_{3}$ and $\sigma_{4}$. One can observe that the fusion outcome of $\sigma_{1}$ and $\sigma_{2}$ will affect the anyonic statistics of $\sigma_{3}$ and $\sigma_{4}$. Assume that the total system is in the even sector (described by a microscopic Pfaffian-type wavefunction with even number of electrons), then braiding $\sigma_{3}$ and $\sigma_{4}$ gives a phase of $\pi/2$ if $\lambda_{1}\ll\lambda_{2}$, but zero phase if $\lambda_{1}\gg\lambda_{2}$. This is because when the interaction favors the $\psi$-type quasihole, the ground state of the total Hamiltonian would contains two unfused $\sigma$-anyons, $\sigma_{3}$ and $\sigma_{4}$, and one $\psi$-anyon that is the result of fusing $\sigma_{1}$ and $\sigma_{2}$. In this case, because of parity conservation, $\sigma_{3}$ and $\sigma_{4}$ must also fuse into a $\psi$, and thus their braiding must gives $\pi/2$. The reverse is true if the interaction favors the $1$-type quasihole instead: the ground state consists of two unfused $\sigma$-anyons and one $1$-anyon, and braiding the two $\sigma$’s gives zero phase. In either case, the braiding phase becomes scalar and completely deterministic just by an introduction of a two-body interaction. This is in contrast to the case where there is no two-body interaction, or when the system contains only unfused $\sigma$-anyons: the ground state is always two-fold degenerate and braiding any two quasiholes results in either a scalar linear combination of $\pi/2$ and 0, or an $\text{SU}(2)$ rotation within the two-fold degenerate ground state due to its non-Abelian properties.
@@
A more complicated situation arises for six quasiholes. Suppose we start with $\sigma_{i}$, $i=1,...,6$. If all six quasiholes are far apart from each other, then the ground state is exactly four-fold degenerate. Now, if we fuse $\sigma_{1}$ and $\sigma_{2}$, then once again the species of anyons present in the ground states are governed by the effect of the specific two-body interaction. This time, the Hilbert space of $\sigma_{3}$, $\sigma_{4}$, $\sigma_{5}$, and $\sigma_{6}$ will always be two-fold degenerate, and in general non-Abelian braiding is expected in both outcomes of $\sigma_{1}$-$\sigma_{2}$ fusion. However, the two cases will have different braiding matrices: if the interaction favors the $1$-anyon ($\psi$-anyon) then $\sigma_{1}$ and $\sigma_{2}$ fuse to a $1$-anyon ($\psi$-anyon), so the braiding of the remaining four quasiholes will follow the odd (even) sector[73],
@@ IV.1 Effects of two-body pseudopotentials
Figure 5: Difference in self-energy of the two types ($E_{1}-E_{\psi}$) with respect to $\hat{V}_{m}^{2bdy}$ for $m=1$ (blue circles), $m=3$ (red stars), $m=5$ (yellow diamons), $m=7$ (purple squares), and m=9 (green crosses). Maximum system size used in calculation has $N_{e}=18$ electrons. A positive energy difference implies that the $\psi$-type quasihole is energetically favorable compared to $1$-type, and vice versa.
Typically, the coefficients $c_{m}$ in Eq.(23) decreases for increasing $m$[67, 25] for realistic interactions. We also see a trend in Fig. 5 that the self-energy difference between the two anyon species decreases with increasing $m$ (for $\hat{V}_{9}^{2bdy}$ the difference is almost zero). It is thus sufficient to consider only the effects of the first few pseudopotentials. Notably, only $\hat{V}_{3}^{2bdy}$ favors the $1$-type anyons, while other pseudopotentials favor the $\psi$-anyons (with $\hat{V}_{1}^{2bdy}$ favoring it most strongly). Understanding the effect of each individual potential is useful for investigating the effects of different tuning parameters, as we present below. Additionally, as several works have suggested the possibility of realizing individual pseudopotentials in experiments[77, 78], the strong opposing effects of $\hat{V}_{1}^{2bdy}$ and $\hat{V}_{3}^{2bdy}$ may become the essential tools for engineering anyon dynamics in the MR system.
@@ IV.2 Effect of quasihole energetics on braiding
Having determined the effects of different pseudopotential terms on favoring a certain type of fused anyon, we can proceed to verify the effect of fusion channel selection on measured braiding statistics by numerical simulation. Here, we employ exact diagonalization (ED) on a system with four quasiholes. Once again, we work on the spherical geometry where position is parametrized by an azimuthal angle $\theta$ and a polar angle $\phi$. We consider a potential pin configuration consisting of three pins, each with a profile described by the Dirac delta function $\delta^{2}(\theta-\theta_{0},\phi-\phi_{0})$ centered at $(\phi_{0},\theta_{0})$. One pin of strength $g_{1}$ fixed at the south pole, and two pins of strength $g_{2}$ located at positions $(\theta,\phi)=(\theta_{0},0)$ and $(\theta_{0},\pi)$ (see Fig. 6a). The two pins are exchanged along a circle of latitude, which is mathematically equivalent to a rotation by $\pi$ around the $z$-axis. Thus, $\hat{L}_{z}$ of the ground state is proportional to the total Berry phase obtained. The exchange phase can then be extracted from this Berry phase in the same manner described in Ref.[56]. The Gaussian profile is chosen for each potential pin because its matrix element in the Landau orbital basis can be computed analytically. All in all, the Hamiltonian used in ED is[69]
$\hat{H}(\theta_{0})$ $=\lambda_{1}\hat{V}_{1}^{2bdy}+\hat{V}+\hat{H}_{\text{pins}}(\theta_{0})$ (24)
@@
[…]
@@
Figure 6: (a) Schematic diagram showing the procedure carried out in numerics. In a system with four MR quasiholes, two quasiholes are each trapped by two potential pins and moved along a circular exchange path on the northern hemisphere, while the other two are fused into a charge-$e/2$ quasihole and fixed by a strong potential pin at the south pole. (b-e) Berry phase as a function of solid angle for a Hamiltonian consisting of a pinning potential for (b-c) 10 electron and (d-e) 11 electrons. In Fig (b) and (d) no two-body interaction is present, while in Fig. (c) and (e) a background interaction of $\hat{V}_{1}^{2bdy}$ is added. In each figure, the exchange phase (in multiple of $2\pi$) is given by the $y$-intercept in the fitting equation indicated on the plot title.
@@
In general, there are two states of interest where the charge-$e/2$ quasihole at the south pole is either the $1$-anyon or the $\psi$-anyon. We first consider the even sector ($N_{e}=8,10,12,...$) and $\lambda_{1}$=0: the ground state in a system with four MR quasiholes is not an exactly two-fold degenerate, as one may expect. This is because the south-pole pin slightly favor the $1$-anyon due to its lower local electron density [^4] This is a unique feature of the Dirac delta potential profile we used. In real experiment, if the one-body potential has a larger profile (e.g. a size scale of several magnetic length) the energy gap between two types of quasiholes could be smaller and the non-Abelian manifold is preserved.. As a result, the other two charge-$e/4$ quasiholes must also fuse into a $1$-type so that the total number of $1$-anyon is even. Since the ground state is now Abelian (it is non-degenerate), the braiding phase of two quasiholes can be related to their topological spins, as well as the topological spin of their fusion product. In this case when two MR quasiholes fuse into a $1$-anyon, one observe a trivial braiding phase between them.
@@
Now we turn on $\hat{V}_{1}^{2bdy}$ by setting $\lambda_{1}>0$. Since the $\psi$-type has lower variational energy with respect to $\hat{V}_{1}^{2bdy}$, as we increase $\lambda_{1}$ there will be level-crossing point beyond which the ground state consists of a $\psi$-type quasihole trapped at the south pole. Now, the free quasiholes must fuse into $\psi$ type so that the total number of $1$-anyon is zero, an even number. Thus, exchanging the free quasiholes result in an exchange phase of $\pi/2$. This change in braiding phase is reflected in the qualitative different between Figs. 6b-c for the even sector. Note that the braiding phase extracted by this method is not exactly quantized at either 00 or $\pi/2$ due to finite size effect[56, 69].
@@
We also note that the effect of $\hat{V}_{1}^{2bdy}$ can also be seen in the odd sector ($N_{e}=7,9,11,...$). In this case, the braiding phase is $\pi/2$ if $\lambda_{1}=0$ and 00 if $\lambda_{1}>0$, the opposite of the results for the even sector. The numerical result in Fig. 6d-e, are however less stable in the odd sector, as evidence in the large jump in the Berry phase data points. This is because unlike the even sector, where the ground state with two potential pins must be of the same type, both of the lowest two ground states in the odd sector contain one $\psi$-anyon and one $1$-anyon. This results in the a much more complicated interplay between anyon self-energy and anyon interaction at finite separation[51]. This complicated anyon dynamics could play a crucial role in designing a suitable braiding scheme, which will be explored further in future work. However, we emphasize that anyon interaction at finite seperation not affect our discussion about the entanglement between fused background anyons and measured braiding statistics, as we can see the qualitative difference between the braiding phases in Figs. 6b and e (close to integer times $\pi$) and those in Figs. 6c and d (close to half integer times $\pi$). In the thermodynamic limits, these quantities should converge to their expected values. This result shows an interesting “long-range entanglement” behavior where what happens to a pair of anyon can affect others that are infinitely far away in a physically measurable manner.

### v3

@@ III Simultaneous selection of fusion channels
Parity conservation places a constrain on the possible outcome of a fusion of two MR quasiholes. To take a simplest example, let us consider the four-quasihole MR state in the even sector. If one pair of quasiholes is fused with a pre-determined outcome, let's say, into a $1$-anyon quasihole, then the fusion out come of the other pair is also known. Now, interesting physics can happen due to the following observation: the fusion outcome of two quasiholes directly affects their mutual statistics[48, 57, 56]. This means that by pre-determining the fusion outcome of one pair of quasiholes, one can directly influence the statistics of other pairs of quasiholes in the system regardless of the distance between them. Thus, parity conservation acts like a long-range entanglement between pairs of MR quasiholes (see Fig. 1).
In this section, we explore the physics resulting from this ``parity-mediated entanglement''. We start this section by a simple analogy between parity conservation in the MR state and typical picture of entanglement in quantum mechanics (i.e. the Bell state). This will motivate discussions on the effect of spontaneous fusion channel selection on braiding measurement, which we discuss subsequently. Numerical evidences for this effect will also be presented in the next section.
@@ III.1 Parity conservation as long-range entanglement
where $S_{i}\uparrow$ denotes the outcome of measuring spin up in the $i$-th electron. Entanglement in this case is evidence in the fact that $S_{1}\uparrow''$ and $S_{2}\uparrow''$ are not independent events.
Coming back to the MR system, suppose we have four MR quasiholes that are separated into two pairs. Each pair, on its own, can fuse into either $1$ or $\psi$. Due to parity conservation, however, the two pairs must both fuse into $1$ or into $\psi$ (here we first take the even sector as an example), so the two possible outcomes can be written as $|\psi,\psi\rangle$ or $|1,1\rangle$. One can then construct a state similar to Eq.(18) as:
@@
Similar to what happens in the Bell state, here the fusion outcome of one pair of anyons, if known, uniquely determines the fusion outcome of the other regardless of their spatial separation. Note that one can similarly construct an entangled state in the odd sector, where there is exactly one $\psi$-anyon and one $1$-anyon:
$\frac{1}{\sqrt{2}}\left(|1,\psi\rangle+|\psi,1\rangle\right)$ (22)
Constructing model ``Bell-type'' states in this manner unveils the mechanism with which parity conservation can affect the physics of a subsystem. Let us next consider a MR system with six quasiholes. In this case, there can be up to three $\psi$-anyons. In the even sector, there can either be zero or two $\psi$-anyons, and in the odd sector there can be either one or three. Without loss of generality, let us take the even sector and consider the following state:
$\frac{1}{2}\left(|\psi,\psi,1\rangle+|\psi,1,\psi\rangle+|1,\psi,\psi\rangle+|1,1,1\rangle\right)$ (23)
which is a possible state with six quasiholes forming three pair, each of which may fuse into either $1$ or $\psi$, but the total number of $\psi$ must be even. We partition the quasiholes into two parts: a ``background'' part consisting of the first pair (whose fusion outcome is denoted by the first symbol in each ket in Eq.(23), and a ``subsystem'' part consisting of the second and third pairs (whose fusion outcomes are denoted by the second and third symbols, respectively). By using the simple calculation of conditional probability as illustrated above, one can verify that the background and subsystem are indeed entangled. For example, suppose we compute the probability of both the second and third pairs fusing into 1:
$P(1_{(2)},1_{(3)})$ $=1/4$ (24)
$P(1_{(2)},1_{(3)}|\psi_{(1)})$ $=0$ (25)
(Here we use the notation $\alpha_{(i)}$ to mean the outcome ``the $i$-th quasihole pair fuses into one $\alpha$-type quasihole''.)
Now, suppose we make a ``measurement'' of the fusion outcome of the first pair of quasiholes. Realistically, this is done by bringing them together and tuning the interaction such that one fusion channel is energetically favoured over the other (the exact details are presented in the next section). In that case, we see that Eq.(23) collapses into either of the following:
$\frac{1}{\sqrt{2}}|1\rangle\otimes\left(|\psi,\psi\rangle+|1,1\rangle\right)$ if outcome is 11 (26)
$\frac{1}{\sqrt{2}}|\psi\rangle\otimes\left(|1,\psi\rangle+|\psi,1\rangle\right)$ if outcome is $\psi$ (27)
Comparing these two with Eqs. (21) and (22), we see that the state describing the ``subsytem'' reduces to one of the two states describing four quasiholes. Interestingly, where as previously Eqs. (21) and (22) belong to topologically distinct Hilbert spaces and never mixes, here both Eqs. (26) and (27) reside within the same Hilbert space. More strikingly is the fact that by fusing a pair of anyons in background and tuning the interaction such that the outcome is $\psi$, we now have a subsystem in Eq. (27), which resembles the odd-sector four-quasihole state, within a system with even parity. We will argue that this ``parity flip'' of a subsystem has a measurable consequence in the anyonic statistics.
We note that all entangled states can only exist within the degenerate nullspace of the parent model Hamiltonian of the MR state, i.e. the three-body interaction $\hat{V}_{3}^{3bdy}$. Under some two-body interaction, the groundstate degeneracy is split, and Eq.(21) collapses into either of each component depending on whether $1$-anyon or $\psi$-anyon are energetically favorable. The exact details on the energetics of different anyon species will be discussed in the following section.
@@ III.2 Effects on braiding statistics
In a FQH system there are two separate ways the fusion outcome of two anyons can be realized and measured: by placing two anyons at the same spatial co-ordinate, effectively ``fusing'' them in real space, or by measuring their mutual statistics via braiding. Assuming that during the braiding process the separation between every pair of $\sigma$-anyons remains very large (compared to their sizes), there exists a qualitative difference between the two methods. Namely, the presence of two-body interaction does not affect the degeneracy of the system during the braiding process, but two-body interaction does split the degeneracy when two $\sigma$-anyons are fused in real space. In the latter case, the fusion channel is chosen energetically depending on the details of two-body interaction (see Fig.4).
@@
which resembles Eq. (1) where $\hat{V}_{3}^{3bdy}$ is the three-body interaction the nullspace of which is the Moore-Read quasihole manifold, $\hat{V}_{\psi}^{\text{model}}$ is some ``model'' two-body interaction that energetically favors $\psi$-anyon over $1$-anyon, and $\hat{V}_{1}^{\text{model}}$ is some other ``model'' two-body interaction that favors $1$-anyon over $\psi$-anyon, and $\hat{V}_{\text{pins}}$ is a collection of some local one-body pinning potentials. In the next section we shall see that indeed $\hat{V}_{1}^{\text{model}}$ and $\hat{V}_{\psi}^{\text{model}}$ exist, but their exact forms are not important for the current discussion [^3] Note that despite the word “model” used, we do not require that $1$-anyon is the zero-energy state of $\hat{V}_{1}^{\text{model}}$, but only that the $1$-anyon has lower energy than $\psi$-anyon with respect to this interaction. The same is true for the $\psi$-anyon and $\hat{V}_{\psi}^{\text{model}}$.
@@
Using this Hamiltonian, one can design an experiment to observe the effect of long-range entanglement as follows. Suppose, in a MR system with four quasiholes, $\sigma_{1}$$\sigma_{2}$$\sigma_{3}$,and $\sigma_{4}$, we pick one pair of quasiholes (say, $\sigma_{1}$ and $\sigma_{2}$) and move their positions to the same spatial point so that they fuse into a charge-$e/2$ quasihole. Afterward, we measure the braiding of $\sigma_{3}$ and $\sigma_{4}$. One can observe that the fusion outcome of $\sigma_{1}$ and $\sigma_{2}$ will affect the anyonic statistics of $\sigma_{3}$ and $\sigma_{4}$. Assume that the total system is in the even sector (described by a microscopic Pfaffian-type wavefunction with even number of electrons), then braiding $\sigma_{3}$ and $\sigma_{4}$ gives a phase of $\pi/2$ if $\lambda_{1}\ll\lambda_{2}$, but zero phase if $\lambda_{1}\gg\lambda_{2}$. This is because when the interaction favors the $\psi$-type quasihole, the ground state of the total Hamiltonian would contains two unfused $\sigma$-anyons, $\sigma_{3}$ and $\sigma_{4}$, and one $\psi$-anyon that is the result of fusing $\sigma_{1}$ and $\sigma_{2}$. In this case, because of parity conservation, $\sigma_{3}$ and $\sigma_{4}$ must also fuse into a $\psi$, and thus their braiding must gives $\pi/2$. The reverse is true if the interaction favors the $1$-type quasihole instead: the ground state consists of two unfused $\sigma$-anyons and one $1$-anyon, and braiding the two $\sigma'$s gives zero phase. In either case, the braiding phase becomes scalar and completely deterministic just by an introduction of a two-body interaction. This is in contrast to the case where there is no two-body interaction, or when the system contains only unfused $\sigma$-anyons: the ground state is always two-fold degenerate and braiding any two quasiholes results in either a scalar linear combination of $\pi/2$ and 0, or an $\text{SU}(2)$ rotation within the two-fold degenerate ground state due to its non-Abelian properties.
@@
A more complicated situation arises for six quasiholes. Suppose we start with $\sigma_{i}$, i=1,,6i=1,...,6. If all six quasiholes are far apart from each other, then the ground state is exactly four-fold degenerate. Now, if we fuse $\sigma_{1}$ and $\sigma_{2}$, then once again the species of anyons present in the ground states are governed by the effect of the specific two-body interaction. This time, the Hilbert space of $\sigma_{3}$, $\sigma_{4}$, $\sigma_{5}$, and $\sigma_{6}$ will always be two-fold degenerate, and in general non-Abelian braiding is expected in both outcomes of $\sigma_{1}$-$\sigma_{2}$ fusion. However, the two cases will have different braiding matrices: if the interaction favors the $1$-anyon ($\psi$-anyon) then $\sigma_{1}$ and $\sigma_{2}$ fuse to a $1$-anyon ($\psi$-anyon), so the braiding of the remaining four quasiholes will follow the odd (even) sector[75],
@@ IV.1 Effects of two-body pseudopotentials
Figure 5: Difference in self-energy of the two types ($E_{1}-E_{\psi}$) with respect to $\hat{V}_{m}^{2bdy}$ for $m=1$ (blue circles), $m=3$ (red stars), $m=5$ (yellow diamons), $m=7$ (purple squares), and m=9 (green crosses). Maximum system size used in calculation has $N_{e}=18$ electrons. A positive energy difference implies that the $\psi$-type quasihole is energetically favorable compared to 11-type, and vice versa[79].
Typically, the coefficients $c_{m}$ in Eq.(29) decreases for increasing mm[67, 25] for realistic interactions. We also see a trend in Fig. 5 that the self-energy difference between the two anyon species decreases with increasing $m$ (for $\hat{V}_{9}^{2bdy}$ the difference is almost zero). It is thus sufficient to consider only the effects of the first few pseudopotentials. Notably, only $\hat{V}_{3}^{2bdy}$ favors the $1$-type anyons, while other pseudopotentials favor the $\psi$-anyons (with $\hat{V}_{1}^{2bdy}$ favoring it most strongly). Understanding the effect of each individual potential is useful for investigating the effects of different tuning parameters, as we present below. Additionally, as several works have suggested the possibility of realizing individual pseudopotentials in experiments[80, 81], the strong opposing effects of $\hat{V}_{1}^{2bdy}$ and $\hat{V}_{3}^{2bdy}$ may become the essential tools for engineering anyon dynamics in the MR system.
Figure 6: (a) Schematic diagram showing the procedure carried out in numerics. In a system with four MR quasiholes, two quasiholes are each trapped by two potential pins and moved along a circular exchange path on the northern hemisphere, while the other two are fused into a charge-$e/2$ quasihole and fixed by a strong potential pin at the south pole. (b-g) The spectrum of Eq.(30) plotted against parameter $\theta_{0}$ for MR systems with: (b) 10 electrons, $\lambda_{1}=0.1$ (c) 10 electrons, $\lambda_{1}=1.0$, (d) 10 electrons, $\lambda_{1}=10.0$, (e) 11 electrons, $\lambda_{1}=0.1$, (f) 11 electrons, $\lambda_{1}=1.0$, (g) 11 electrons, $\lambda_{1}=10.0$[79].
@@ IV.2 Effect of quasihole energetics on braiding
Having determined the effects of different pseudopotential terms on favoring a certain type of fused anyon, we can proceed to verify the effect of fusion channel selection on measured braiding statistics by numerical simulation. Here, we employ exact diagonalization (ED) on a system with four quasiholes. Once again, we work on the spherical geometry where position is parametrized by an azimuthal angle $\theta$ and a polar angle $\phi$. We consider a potential pin configuration consisting of three pins, each with a profile described by the Dirac delta function $\delta^{2}(\theta-\theta_{0},\phi-\phi_{0})$ centered at $(\phi_{0},\theta_{0})$. One pin of strength $g_{1}$ fixed at the south pole, and two pins of strength $g_{2}$ located at positions $(\theta,\phi)=(\theta_{0},\phi_{0}(t))$ and $(\theta_{0},\phi_{0}(t)+\pi)$ (see Fig. 7a). The polar position $\phi_{0}(t)$ is then varied between 0 and $\pi$ such that two pins are exchanged along a circle of latitude, which is mathematically equivalent to a rotation by $\pi$ around the $z$-axis. Thus, $\hat{L}_{z}$ of the ground state is proportional to the total Berry phase obtained. The exchange phase can then be extracted from this Berry phase in the same manner described in Ref.[56]. The Gaussian profile is chosen for each potential pin because its matrix element in the Landau orbital basis can be computed analytically. All in all, the Hamiltonian used in ED is[69]
$\hat{H}(\theta_{0})$ $=\lambda_{0}\hat{V}_{3}^{3bdy}+\lambda_{1}\hat{V}_{1}^{2bdy}+\hat{H}_{\text{pins}}(\theta_{0})$ (30)
@@
$\hat{V}_{g}(\theta_{0},\phi_{0})$ $=\int d\Omega\hat{\rho}(\theta,\phi)\delta^{2}(\theta-\theta_{0},\phi-\phi_{0})$ (32)
@@
In our numerics, we take the limit $\lambda_{0}\to\infty$ so that all low-lying states reside within the MR CHS spanned by the Jack polynomials. Even then, there are two possible tuning parameters: $\lambda_{1}$, which is the strength of $\hat{V}_{1}^{2bdy}$, and $g_{1}$, which is the strength of the potential pin. We fix $g_{1}=1$ and investigate the effect of tuning $\lambda_{1}$ on the spectrum. Based on the previous analysis, we know that $\hat{H}_{\text{pins}}$ slightly favours the $1$-anyon energetically while $\hat{V}_{1}$ favours the $\psi$-anyon, and thus we expect a crossing between the lowest two energy levels when the charge-$e/2$ anyon pinned at the south pole transmutes from $1$-type to $\psi$-type. However it is not so clear in numerics, as seen in the spectrum in Figs. 6b-g. This is because the spectrum is complicated by the presence of the two charge-$e/4$, which interacts at finite separation. The behavior in the even sector ($N_{e}=10$ electrons, Figs. 6b-d) is relatively simple: at $\theta_{0}=0$, the ground state is either two $1$-anyons, one at each pole, for small $\lambda_{1}$, or two $\psi$-anyons, also one at each pole, for large $\lambda_{1}$. As $\theta_{0}$ is increased (at fixed $\lambda_{1}$), the charge-$e/2$ quasihole pinned at the south pole remains a fixed type, and the interaction between the two charge-$e/4$ results in the fluctuation in energy of the ground state (the quantitative behavior of this effective anyon interaction is known from Ref. [51]).
On the other hand, the odd sector ($N_{e}=11$ electrons, Fig. 6e-g is more complicated since at $\theta_{0}$=0 the ground state is two-fold degenerate: in the system there is one $1$-anyon and one $\psi$-anyon, the former of which may be pinned at either the north pole or the south pole, and the latter will be pinned at the opposite pole. Because of this, when $\theta_{0}$ are increased, the charge-$e/2$ quasihole at the south pole may spontaneously transmute between the two anyon types, resulting in level crossings between the two lowest energies. However, since these level crossings are a results of balancing between interaction between the two charge-$e/4$ anyons and the self-energy of the charge-$e/2$ anyon, they are not expected to occur at large $\theta_{0}$, where the charge-$e/4$ quasiholes are separated enough from each other such that they are non-interacting. Indeed, in Figs. 6e-g we see that the crossings only occur at relatively small values of $\theta_{0}$ (less than 0.5 radian).
Figure 7: Berry phase as a function of solid angle $\Omega=2\pi(1-\cos\theta_{0})$ in the braiding scheme using Eq. 30 described in the main text. A linear fit is performed on a few right-most data points ($\Omega\sim 0.5$), from which the exchange statistics in multiple of $2\pi$ can be taken as the $y$-intercept (shown in the title of each respective plot) (a) 8 electrons, $\lambda_{1}=0$ (b) 9 electrons, $\lambda_{1}=0$ (c) 10 electrons, $\lambda_{1}=0$ (d) 11 electrons, $\lambda_{1}=0$ (e) 8 electrons, $\lambda_{1}>0$ (f) 9 electrons, $\lambda_{1}>0$ (g) 10 electrons, $\lambda_{1}>0$ (h) 11 electrons, $\lambda_{1}>0$[79].
@@
The Berry phase resulting from this braiding scheme is shown in Fig. 7 for systems with $N_{e}=8,9,10$, and $11$ electrons. We first consider the even sector ($N_{e}=8$ and $10$) and $\lambda_{1}$=0: the ground state in a system with four MR quasiholes is not an exactly two-fold degenerate, as one may expect. This is because the south-pole pin slightly favor the $1$-anyon due to its lower local electron density [^4] This is a unique feature of the Dirac delta potential profile we used. In real experiment, if the one-body potential has a larger profile (e.g. a size scale of several magnetic length) the energy gap between two types of quasiholes could be smaller and the non-Abelian manifold is preserved.. As a result, the other two charge-$e/4$ quasiholes must also fuse into a $1$-type so that the total number of $1$-anyon is even. Since the ground state is now Abelian (it is non-degenerate), the braiding phase of two quasiholes can be related to their topological spins, as well as the topological spin of their fusion product. In this case when two MR quasiholes fuse into a $1$-anyon, one observe a trivial braiding phase between them.
@@
Now we turn on $\hat{V}_{1}^{2bdy}$ by setting $\lambda_{1}>0$. Since the $\psi$-type has lower variational energy with respect to $\hat{V}_{1}^{2bdy}$, as we increase $\lambda_{1}$ there will be level-crossing point beyond which the ground state consists of a $\psi$-type quasihole trapped at the south pole. Now, the free quasiholes must fuse into $\psi$ type so that the total number of $1$-anyon is zero, an even number. Thus, exchanging the free quasiholes result in an exchange phase of $\pi/2$. This change in braiding phase is reflected in the qualitative difference between Figs. 7a and e for 8 electrons and between Figs. 7c and g for 10 electrons. (In this numerics, we choose $\lambda_{1}=1$ for the even sector so that the spectrum is qualitatively different from the case when $\lambda_{1}=0$ , see Fig. 6c.) Note that the braiding phase extracted by this method is not exactly quantized at either 0 or $\pi/2$ due to finite size effect[56, 69]. However, one can see the qualitative difference in the braiding phase $\gamma_{br}$ between $\gamma_{br}\sim 0$ at $\lambda_{1}=0$ and $\lambda_{br}\sim\pi/2$ at $\lambda_{1}>0$
@@
We also note that the qualitative effect of $\hat{V}_{1}^{2bdy}$ can also be seen in the odd sector ($N_{e}=7,9,11,$...), though the quantitative result depends on the complex interaction among all three types of anyons and remains to be studied further. In this case, the braiding phase is $\pi/2$ if $\lambda_{1}=0$ and 0 if $\lambda_{1}>0$, the opposite of the results for the even sector. The numerical results, shown in Fig. 7b and f for 9 electrons and Fig. 7d and h for 11 electrons, are less stable compared to their even-sector counterparts, as evidence in the large jump in the Berry phase data points. These jumps are a direct consequence of the level crossings between the lowest two energy states, as discussed above and seen in Fig. 6e-g. This complicated anyon dynamics could play a crucial role in designing a suitable braiding scheme, which will be explored further in future work. However, we emphasize that anyon interaction at finite seperation not affect our discussion about the entanglement between fused background anyons and measured braiding statistics, as we can see the qualitative difference between the braiding phases in Figs. 7b and e (close to integer times $\pi$) and those in Figs. 7c and d (close to half integer times $\pi$). In the thermodynamic limits, these quantities should converge to their expected values. This result shows an interesting ``long-range entanglement" behavior where what happens to a pair of anyon can affect others that are infinitely far away in a physically measurable manner.